OpenStudy (anonymous):

10 is the geometric mean of 4 and what number?

OpenStudy (anonymous):

\[\sqrt{4*x}=10\] 4*x = 100 x=25

OpenStudy (anonymous):

geometric mean \[G=\sqrt[n]{x_{1}x_{2}x_{3}...x_{n}}\]

Directrix (directrix):

When a positive value is repeated in either the means or extremes position of a proportion, that value is referred to as a geometric mean (or mean proportional) between the other two values. 4 is to 10 as 10 is to x where x is the number for which 10 is the geometric mean of 4 and x. 4/10 = 10/x 4x = 100 x = 25